Ch 11: Exploring Finance Visualizations -Short-Term versus Long-Term Cash Flows the discounted value when the interest rate (1. What is the percentage change in the PV of $100 due in 1 year when the interest rate changes from 5% to 10%? (Move the sli4. If the interest rate is less than 5%, then the PVs for both the one-year and 20-year investments: a. Decrease because theHi this is all one question.
Answer #1

Answer 1)

Time = 1 year

Amount = $100

PV when r = 5%

PV = 100 / (1 + 5%)^1

PV = $95.24

PV when r = 10%

PV = 100 / (1 + 10%)^1

PV = $ 90.91

Hence, % change = (90.91 - 95.24) * 100 / 95.24

Change = - 4.55%

Hence option C (Decrease by 4.5%)

------------------------------------------------------

Answer 2)

Time = 20 year

Amount = $100

PV when r = 5%

PV = 100 / (1 + 5%)^20

PV = $ 37.69

PV when r = 10%

PV = 100 / (1 + 10%)^20

PV = $ 14.86

Hence, % change = (14.86 - 37.69) * 100 / 37.69

Change = - 60.56%

Hence option A (Decrease by 60%)

------------------------------------------------------

Answer 3)

Correct option is Option C. We can see that time period increases the denominator value and hence PV is more small if interest rate changes in long term compared to short term

------------------------------------------------------

Answer 4)

Correct option is Option B. PV for both will increase as the denominator in formula will be smaller and hence higher PV.